Valuing BondsEven though most corporate bonds in the United States make coupon payments semiannually, bonds issued elsewhere often have annual coupon payments. Suppose a German company issues a bond with a par value of €1,000, 15 years to maturity, and a coupon rate of 5.1 percent paid annually. If the yield to maturity is 4.3 percent, what is the current price of the bond?
Answer #1

Annual coupon=1000*5.1%=51

Hence price of bond=Annual coupon*Present value of annuity factor(4.3%,15)+1000*Present value of discounting factor(4.3%,15)

=51*10.88874113+1000*0.531784131

which is equal to

=1087.11(Approx)

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=51[1-(1.043)^-15]/0.043

=51*10.88874113

2.Present value of discounting factor=1000/1.043^15

=1000*0.531784131